From GlueXWiki
Jump to: navigation, search

April 17, 2014 FDC meeting


  1. FDC testing status (Lubomir)
    1. Final scope results
    2. Beni's results with DAQ
  2. Engineering (Bill, Dave)
    1. Gas system status [1],[2]
    2. Experimental Safety Assessment Document HallD_ESAD
    3. Other
  3. Electronics (Chris, Nick)
  4. Other


Participants: Eugene, Fernando, Dave, Nick, Simon, and Lubomir.

FDC testing status

- Final results from the scope test (done by Anatoly) of all the channels with FDC in its final position are in the spreadsheet linked above. About 50 strip channels (0.4% of all the strips) are identified as having some kind of problem, almost all of them related to the conductive tape issue. As compared to the scope tests done after the first "final" insertion almost nothing changed, i.e. the re-insertion didn't create additional problems.

- Comparing with the Beni's DAQ results: all the problematic channels from the scope test can be identified in the Beni's histograms as deviating from the neighboring channels, except several channels that are probably dominated by a noise. In addition to the 50 bad channels from the scope test, there are about 30 new bad channels in the Beni's results. Lubomir looked with a scope into 10 of these channels and didn't find any problem. Fernando and Chris did similar test with a scope and concluded for most of the channels the signals at the input of the fADCs are good. However, according the Beni the problematic channels move with the cable. Therefore Fernando suspects it might be the front panel on the fADC that prevents the cable connector from completely going in the fADC connector. They will investigate this.


- Last week the FDC was moved downstream by about 5mm (info from Josh) to make space for the CDC, without releasing the FDC cables on the face of the magnet. According to Casey they just looked to make sure there's no tension on the cables. Then stoppers were installed on both FDC and CDC side on the rail. On Tuesday survey of the FDC in its final position was done. Important is that they were able to see all the 16 reflectors installed on the periphery of the packages. Before that the CDC was surveyed: it is off by 7mm downstream with respect to the nominal position (Eugene).

- The gas system is next on the list, the OSP for the pressure testing will be signed soon and beginning of next week the test will be done. Dave is working on the procedure (files attached) for testing and flushing the whole gas system which is suppose to happen by the end of next week. Yesterday the mass flow controller that was used temporary for the FDC was needed to complete the gas system and was replaced with a rotameter that somehow consumes much more gas. Fernando: flushing with alcohol will require also OSP. That's why the flushing will be done for now without alcohol, later it will be added.

- Elton is working on the ESAD document and we are responsible for the FDC and the gas sections there. The first version is attached above (pages 14-15). Fernando: the HV and LV hazards will be described in the electronics section, so we just have to refer to that section.

- At the beginning of the week the Fluorinert level was very low and we put the last gallon we had, in the chiller and even after that the level is much below the "low" mark. We suspect some leakage and will need the Fluorinert sniffer again to check at least the outside connections. Dave ordered another container (20kg) with Fluorinert.


- Fernando: the LV supplies are now regulating the sense voltage, i.e. the voltage on the patch panels. The numbering convention for the electronics channels should be followed, i.e. first channels is 0.

- The FDC ground is now connected directly to the clean ground while the CDC ground is not connected. There's no ground connection for the mesh: it might be difficult but not impossible, to connect a ground cable on the side of the mesh. Soon the ground panel on the face of the magnet will be attached to the clean ground and the FDC, CDC, and the mesh grounds will be attached to that panel.

CDC update

- Due to the big overlap in equipment, procedures, software (gas system, pre-amps, fADC, HV, LV modules, DAQ, testing software) with the CDC and the experts present at this meeting, Fernando proposed and everybody agreed that this meeting is the right place to give updates also on the CDC activities.

- Nick and two techs finished the connections of the signal cables to the CDC. Now in principle one can apply LV on the pre-amps and Beni can look at the pedestal widths with the mini-DAQ to check the cable connectivity. Before that the cooling needs to be finished and make temperature measurements on the pre-amps. After the test of the gas system, in about 10-14 days, one can start flushing the CDC.